OpenStudy (anonymous):

r/6=5 2/3 solve A 34 B 30 2/3 C 20 D 17/18

OpenStudy (leader):

5 2/3=17/3

OpenStudy (leader):

\[\frac{ r }{ 6 }=\frac{ 17 }{3 }\]

OpenStudy (leader):

So 3r=102 r=34

OpenStudy (leader):

@Mr.BigPersnolty